OpenStudy (anonymous):

the line (y-2) = k(x+1) passes through the points L(3,3) and M(7,4) find K

OpenStudy (anonymous):

K is the slope. You get the slope by applying this formula to the two points you've already provided: \[\frac{ y-y_1 }{ x-x_1 }\]. So you have: \[\frac{ 4-3 }{ 7-3 }\] Now you simplify. \[\frac{ 1 }{ 4 }\]=K

OpenStudy (anonymous):

ok thanks soo much :)

OpenStudy (anonymous):

You're welcome